Find an equation of the tangent line to the graph of y = x2 + 3x at the point (-1, -2)
show clear work plz
Answer #1

We have to find an equation of the tangent line to the graph of

y=x^{2}+3x

at the point (-1,-2).

Now, the equation of the tangent line at the point (1,-2) is

y-(-2)=\left | \frac{dy}{dx} \right |_{-1,-2}(x-(-1))

Now, from the equation of the graph, we find the derivative.

Now,

y=x^{2}+3x

or, \frac{dy}{dx}=2x+3

So, at the point (-1,-2), the value of the derivative becomes

\frac{dy}{dx}_{(-1,-2)}=2(-1)+3=1

So, the slope of the straight line is 1.

This means, the equation of the tangent line is

y-(-2)=1[x-(-1)]

or, y+2=x+1

or, x-y=1

Thus, the equation is x-y=1.

The answer is

The equation of the tangent line to the graph of

y=x^{2}+3x

at the point (-1,-2) is given by the equation

x-y=1.